# The neutrino behavior in the Einstein-Hilbert-Bumblebee gravity around global monopole field

### v1

@@ Abstract
In this work, we study the phase of neutrino oscillation propagating along radial and non-radial paths in the Einstein-Hilbert-Bumblebee (EHB) gravity around global monopole field, by using the quantummechanical treatment, the expression of the corrected neutrino of probability of oscillation is obtained. Our results show that neutrino oscillation in the EHB gravity around global monopole field is different from that in Schwarzschild black hole, the Lorentz-violating parameter $a$ mainly affects the peak of the oscillation probability, and the global monopole $\overline{\mu}$ and the lightest neutrino mass both affect the frequency of the oscillations of probabilities, which means that studying the neutrino oscillation phenomenon in curved spacetime may not only become a new technology for probing the properties of compact celestial bodies, but also help deepen our understanding of neutrinos.
@@ I Introduction
In this work, we study the neutrino behavior in the EHB gravity around global monopole field. In this line of thought, considering the existence of lensing around a massive astrophysical object, the oscillation probability at a particular point is calculated by analysing all possible trajectories of neutrinos which get focused at that point due to the lensing phenomenon retChakrabarty . General relativity and the standard model of particle physics are two important theoretical cornerstones of modern physics, and unifying these two theories would give us a deeper understanding of nature. The theory of quantum gravity is an important product in the process of seeking the grand unification of theories. When it comes to the numerous modified gravitational theories, a possible extension of general relativity is the theory of gravity with Lorentz symmetry breaking retJacobson ; retHorava , which is related to the view that the quantum gravity signals are likely to be detected at low energy scales retCASA , and another compelling point is that this work retCASA opens up the scope of the static and spherically symmetric black hole solutions in bumblebee gravity, subsequently this research is widely applied to various scenarios, including the scalar quasibound states in Einstein-Maxwell-Bumblebee black holes retSenjaya:2025cgk , the field perturbations and observables in a charged black hole within bumblebee gravity retSingh:2025fuv , the energy extraction from a rotating black hole via magnetic reconnection retYuChih:2025hsg , the electrodynamic black holes in Bumblebee gravity with Lorentz Violation and shadow phenomenology retSekhmani:2025gvv , the applications of gauss-bonnet theorem retOvgun:2018ran , the ergosphere dynamics and rotational energy extraction Hassanabadi:2025ect and the quasinormal modes retDeng:2025uvp . As a pillar of modern physics, Lorentz symmetry may not be valid at the Planck scale retMattingly , therefore testing possible Lorentz symmetry breaking signatures in the gravitational sector is becoming a hot research direction, especially in terms of how this mechanism affects the thermodynamics of black holes retFilho:2021rin ; retAnacleto:2018wlj ; retFilho:2020cgk ; retFilho:2023ydb . The bumblebee gravity is one of the Lorentz violation gravitational theories, thus it has been received a lot of attention retShi:2025rfq ; retShi:2025plr . On the other side, spontaneously broken global symmetries give rise to a divergence of the global monopole energy in terms of the monopole case, a common approach dealing with this problem is to introduce the self gravity of global monopole, and the effect of global monopole on Schwarzschild spacetime has been studied in retDadhich .
@@
Based on the above considerations, in this manuscript we focus on the neutrino behavior in the EHB gravity around global monopole field aiming to study the effect of the weak gravitational lensing on the neutrino oscillation. The article is organised as follows: a brief review of Schwarzschild-like black hole with a topological defect in bumblebee gravity and the theoretical background and assumptions of neutrino oscillation are provided in sections II and III, respectively. In section IV, we discuss the phase of neutrino oscillation in the EHB gravity around global monopole field and in section V the neutrino oscillation probability in the EHB gravity around global monopole field is studied, the numerical investigation of two flavor case is discussed in section VI, finally we summarize and comment on our results in section VII.
@@ II Schwarzschild-like black hole with a topological defect in bumblebee gravity
where $M$, $a$ and $\overline{\mu}$ represent the usual geometrical mass, the bumblebee parameter and the global monopole term respectively. By introducing the coordinate transformation, the event horizon of the black hole locates at $g_{tt}(r_{h})=0$ so that $r_{h}=2M$, which does not depend on $a$ and $\overline{\mu}$, this implies that bumblebee field and global monopole field has no effect on event horizon retGullu:2020qzu . On the other hand, there exists a physical singularity at $r=0$ because of the divergence of the Kretschmann invariant. Moreover, the Kretschmann scalar is different than the Schwarzschild case, which means that it is impossible to recover Eq. (1) to Schwarzschild spacetime through a coordinate transformation, because this formalism is a real solution containing the global monopole term and the Lorentz-violating corrections. Therefore, in this paper we only consider the gravitational field in the range of radial coordinates $r\in\left(2M,\infty\right)$.
@@ III Theoretical background and assumptions of neutrino oscillation
where $i=1,2,3$ and $U_{\alpha i}^{\ast}$ is a 3x3 neutrino flavor mixing unitary matrix retMaki . Here we assume that the neutrino wave-function have a plane-wave wavefunction retSwami and it propagates in vacuum from source $S(t_{S},\boldsymbol{x}_{S})$ to detector $D(t_{D},\boldsymbol{x}_{D})$, thus during the process of propagation, the phase accumulation of $\lvert\ \upsilon_{i}\rangle$ can be described by $\left\lvert\upsilon_{i}\left(t_{D},\boldsymbol{x}_{D}\right)\right\rangle=\exp\left(-i\varphi_{i}\right)\left\lvert\upsilon_{i}\left(t_{S},\boldsymbol{x}_{S}\right)\right\rangle$, where $\varphi_{i}$ is the phase of oscillation. In flat spacetime, the phase is given by
@@
$\mathcal{P}_{\alpha\beta}\equiv\left\lvert\left\langle v_{\beta}\lvert v_{\alpha}(t_{D},\boldsymbol{x}_{D})\right\rangle\right\lvert^{2}=\sum_{i,j}U_{\beta i}U_{\beta j}^{*}U_{\alpha j}U_{\alpha i}^{*}exp[-i(\varphi_{i}-\varphi_{j})],$ (6)
which means that when the initial and final flavor eigenstates are determined, the probability of oscillation $\mathcal{P}_{\alpha\beta}$ depends only on the phase difference $\Delta\varphi_{ij}=\varphi_{i}-\varphi_{j}$, in the following section, we will focus on calculating the phase accumulation of neutrinos during their propagation in curved spacetime. Inspired by the work retBilenky:1978nj , we assume that all the mass eigenstates in a flavour eigenstate initially produced at the source $S(t_{s},x_{s})$ have equal energy or momentum.
@@ IV Phase of neutrino oscillation in the EHB gravity around global monopole field
Now let’s calculate the phase of the relativistic neutrino oscillation probability described by the metric (Eq. (1)), for simplicity, the motion of neutrinos can be chosen to be confined on $\left(\theta=\frac{\pi}{2}\right)$ plane, thus the components of canonical momenta $p_{\mu}^{\left(h\right)}$ become
@@
$p_{r}^{\left(h\right)}=m_{h}g_{rr}\frac{dr}{ds}=m_{h}\frac{1+\ell}{A}\frac{dr}{ds},$
@@
$\varphi_{h}\approx\pm\int_{r_{s}}^{r_{D}}\frac{E_{h}}{A}\sqrt{1+a\frac{Am_{h}^{2}}{2E_{h}^{2}}}dr=\sqrt{1+a\frac{m_{h}^{2}}{2E_{0}^{2}}}(r_{D}-r_{s}),$ (10)
@@ IV.2 Non-radial propagating neutrino phase in a null trajectory($d\phi\neq 0$)
The phase of neutrinos travelling along a null trajectory in terms of the non-radial case can be written as
@@
where b is the impact parameter. We can see that this equation will recover the phase in flat spacetime when $\overline{\mu}=a=b=0$, while for the case of $\ell=b=0$, the phase is not only affected by the geometry of spacetime, but also by the trajectory of the neutrino from the propagation source to the detector. Therefore, now we need to consider two cases to discuss the phase:
(a) Neutrinos are produced at the source $S$, then propagate outwards non-radially until they reach the detector $D$. Since it is difficult to integrate Eq. (11) by replacing the metric coefficients, we rewrite it by using weak field approximation $M\ \ll\ r$ to get
@@
obviously, the above formula will reduce to that of the Schwarzschild case when $\overline{\mu}=a=0$ retChakrabarty:2023kld .
(b) Neutrinos are produced at the source $S$, propagate towards the gravitational lens, and finally reaches the detector $D$, here $r=r_{0}$ represents the closest distance to the lens during the path. Meanwhile we apply the weak field approximation $M\ \ll\ r$, thus the phase transforms into
@@
$r_{0}\approx b\sqrt{1-\mu}-\frac{M}{(1-\mu)}.$ (16)
@@ V Neutrino oscillation probability in the EHB gravity around global monopole field
As depicted in Fig. 1, neutrinos are produced in a common source and propagate outward along different paths. Due to the presence of a strong gravitational field (Schwarzschild-like black hole), the trajectories of some neutrinos deviate and converge at a common detection point $D$, resulting in the interference phenomenon, which is similar to how a convex lens (equivalent to a black hole) can converge the optical rays of the sun (equivalent to the propagation of neutrinos). Considering the fact that neutrinos traveling through different paths from S to D, so the neutrino flavour eigenstate need to be expressed as retSwami:2021wbf
$|v_{\alpha}(t_{D},\boldsymbol{x}_{D})\rangle=N\sum_{i}U_{\alpha i}^{*}\sum_{p}\exp(-i\varphi_{i}^{p})|v_{i}(t_{s},\boldsymbol{x}_{s})\rangle,$ (18)
where $\varphi_{i}^{p}$ is given by Eq.(17) and $N$ is a normalization factor. Considering the transition from an initial flavor eigenstate $\alpha$ at the source $S$ to a final flavor eigenstate $\beta$ at the detector $D$, the probability of oscillation can be expressed as
$\mathcal{P}_{\alpha\beta}^{lens}\equiv\left|\left\langle v_{\beta}|v_{\alpha}\left(t_{D},\boldsymbol{x}_{D}\right)\right\rangle\right|^{2}=|N|^{2}\sum_{i,j}U_{\beta i}U_{\beta j}^{*}U_{\alpha j}U_{\alpha i}^{*}\sum_{p,q}\exp\left(-i\Delta\varphi_{ij}^{pq}\right),$ (19)
@@
$|N|^{2}=\left(\sum_{i}\left|U_{\alpha i}\right|^{2}\sum_{p,q}\exp\left(-i\Delta\varphi_{ii}^{pq}\right)\right)^{-1}.$ (20)
As mentioned before, when the initial and final flavor eigenstates are determined, the probability of oscillation $\mathcal{P}_{\alpha\beta}^{lens}$ depends on the phase difference, thus in order to facilitate analysis, $\Delta\varphi_{ij}^{pq}$ can be represented by Eq.(17) as
@@
$A_{pq}=\frac{\sqrt{1+\ell}}{{2E}_{0}}\left(r_{D}+r_{s}\right)\left[1+\frac{2M}{r_{D}+r_{s}}-\frac{\sum b_{pq}^{2}}{4r_{D}r_{s}}\right],$ (22)
$B_{ij}=-\frac{\sqrt{1+\ell}}{{8E}_{0}}\sum m_{ij}^{2}\left(\frac{1}{r_{D}}+\frac{1}{r_{s}}\right),$
@@
Observing this form, we can see that if $p\neq q$, it is possible for the probability of oscillation $\mathcal{P}_{\alpha\beta}^{lens}$ to include $\sum m_{ij}^{2}$ term, therefore, in principle, the observation of neutrinos in the lens could potentially detect the information about their absolute mass retChakrabarty .
@@
$P_{e\mu}^{lens}=|N|^{2}\sin^{2}2\gamma\left\{\frac{1}{2}\cos(\Delta b_{12}^{2}B_{11})+\frac{1}{2}\cos(\Delta b_{12}^{2}B_{22})+\sin^{2}\left(\frac{\Delta m_{21}^{2}A_{11}}{2}\right)\right$. (23)
$+\sin^{2}\left(\frac{\Delta m_{21}^{2}A_{22}}{2}\right)-\cos(\Delta m_{21}^{2}A_{12})\cos(\Delta b_{12}^{2}B_{12})\},$
where $|N|^{2}=\frac{1}{2}\left[1+\cos^{2}\gamma\cos\left(\Delta b_{12}^{2}B_{11}\right)+\sin^{2}\gamma\cos\left(\Delta b_{12}^{2}B_{22}\right)\right]^{-1}$ and $\gamma$ is the mixing angle retEsteban:2018azc .
Next, let’s make an analysis for Eq.(23):
(i) Since $\Delta b^{2}$ belongs to an even function, $\mathcal{P}_{\alpha\beta}^{lens}$ does not change under the interchange of $b_{p,q}$,
(ii) $\mathcal{P}_{\alpha\beta}^{lens}$ depends on the neutrino mass ordering and for $\Delta m^{2}>0$ and $\Delta m^{2}<0$, it results in different results.
(iii) the term $B_{12}$ explicitly contains the sum of squared neutrino masses $\sum m_{ij}^{2}$ .
VI Numerical investigation: two flavor case
We prefer to express the path-dependent impact parameter $b_{p}$ in terms of the geometric quantities of the system. As shown in Fig. 1, it shows the description of lensing phenomena, in which the physical distance from the source to the lens is $r_{S}$ and the physical distance from the lens to the detector is $r_{D}$. For convenience, a new coordinate system $(x^{\prime},y^{\prime})$ is introduced by rotating the $(x,y)$ coordinate system through an angle $\Theta$ such that $x^{\prime}=x\cos{\Theta}+y\sin{\Theta}$ and $y^{\prime}=y\cos{\Theta}-x\sin{\Theta}$. Inspired by the work retCASA , the angle of deflection $\delta$ of neutrinos in the $(x^{\prime},y^{\prime})$ coordinate system can be expressed as

### v2

@@ Abstract
In this work, we study the neutrino oscillation phase propagating along radial and non-radial paths in the Einstein-Hilbert-Bumblebee (EHB) gravity around global monopole field, by using the quantummechanical treatment, the expression of the corrected neutrino oscillation probability is obtained. Our results show that neutrino oscillation in the EHB gravity around global monopole field is different from that in Schwarzschild black hole, the Lorentz-violating parameter $a$ mainly affects the peak of the oscillation probability, and the global monopole $\overline{\mu}$ and the lightest neutrino mass both affect the frequency of the oscillations of probabilities, which means that studying the neutrino oscillation phenomenon in curved spacetime may not only become a new technology for probing the properties of compact celestial bodies, but also help deepen our understanding of neutrinos.
@@ I Introduction
In this work, we study the neutrino behavior in the EHB gravity around global monopole field. In this line of thought, considering the existence of lensing around a massive astrophysical object, the oscillation probability at a particular point is calculated by analysing all possible trajectories of neutrinos which get focused at that point due to the lensing phenomenon retChakrabarty . General relativity and the standard model of particle physics are two important theoretical cornerstones of modern physics, and unifying these two theories would give us a deeper understanding of nature. The theory of quantum gravity is an important product in the process of seeking the grand unification of theories. When it comes to the numerous modified gravitational theories, a possible extension of general relativity is the theory of gravity with Lorentz symmetry breaking retJacobson ; retHorava , which is related to the view that the quantum gravity signals are likely to be detected at low energy scales retCASA , and another compelling point is that this work retCASA opens up the scope of the static and spherically symmetric black hole solutions in bumblebee gravity, subsequently this research is widely applied to various scenarios, including the scalar quasibound states in Einstein-Maxwell-Bumblebee black holes retSenjaya:2025cgk , the field perturbations and observables in a charged black hole within bumblebee gravity retSingh:2025fuv , the energy extraction from a rotating black hole via magnetic reconnection retYuChih:2025hsg , the electrodynamic black holes in Bumblebee gravity with Lorentz Violation and shadow phenomenology retSekhmani:2025gvv , the applications of gauss-bonnet theorem retOvgun:2018ran , the ergosphere dynamics and rotational energy extraction Hassanabadi:2025ect and the quasinormal modes retDeng:2025uvp . As a pillar of modern physics, Lorentz symmetry may not be valid at the Planck scale retMattingly , therefore testing possible Lorentz symmetry breaking signatures in the gravitational sector is becoming a hot research direction, especially in terms of how this mechanism affects the thermodynamics of black holes retFilho:2021rin ; retAnacleto:2018wlj ; retFilho:2020cgk ; retFilho:2023ydb . The bumblebee gravity is one of this Lorentz violation gravitational theories, thus it has been received a lot of attention retShi:2025rfq ; retShi:2025plr . On the other side, spontaneously broken global symmetries give rise to a divergence of the global monopole energy in terms of the monopole case, a common approach dealing with this problem is to introduce the self gravity of global monopole. Global monopoles are point-like topological defects, which are assume that emerge during symmetry-breaking phase transitions in the early universe retBarriola:1989hx , especially involving scalar fields, they undergo a process from the global O(3) symmetry into U(1) symmetry during phase transitions in the early universe retVilenkin:1981kz ; retVilenkin:1984ib . Recently, various scenarios are discussed in terms of global monopoles, including the phantom global monopole retNooriGashti:2025vyg , the ray geodesic and wave propagation in lorentz-violating wormholes retAhmed:2025yyb , the global monopole induced wormholes in power-law gravity retYousaf:2025jqx , the black holes surrounded by a Murnaghan-fluid scalar gas retSekhmani:2025yjf and the effect of global monopole on Schwarzschild spacetime retDadhich .
@@
Based on the above considerations, in this manuscript we focus on the neutrino behavior in the EHB gravity around global monopole field aiming to study the effect of the weak gravitational lensing on the neutrino oscillation. The article is organised as follows: a brief review of Schwarzschild-like black hole with a topological defect in bumblebee gravity and the theoretical background and assumptions of neutrino oscillation are provided in sections II and III, respectively. In section IV, we discuss the neutrino oscillation phase in the EHB gravity around global monopole field and in section V the neutrino oscillation probability in the EHB gravity around global monopole field is studied, the numerical results of two flavor case is discussed in section VI, finally we summarize and comment on our results in section VII.
@@ II Schwarzschild-like black hole with a topological defect in bumblebee gravity
where $M$, $a$ and $\overline{\mu}$ represent the usual geometrical mass, the Lorentz-violating parameter and the global monopole term respectively. By introducing the coordinate transformation, the event horizon of the black hole locates at $g_{tt}(r_{h})=0$ so that $r_{h}=2M$, which does not depend on $a$ and $\overline{\mu}$, this implies that bumblebee field and global monopole field has no effect on event horizon retGullu:2020qzu . On the other hand, there exists a physical singularity at $r=0$ because of the divergence of the Kretschmann invariant. Moreover, the Kretschmann scalar is different than the Schwarzschild case, which means that it is impossible to recover Eq. (1) to Schwarzschild spacetime through a coordinate transformation, because this formalism is a real solution containing the global monopole term and the Lorentz-violating corrections. Therefore, in this paper we only consider the gravitational field in the range of radial coordinates $r\in\left(2M,\infty\right)$.
@@ III Theoretical background and assumptions of neutrino oscillation
where $i=1,2,3$ and $U_{\alpha i}^{\ast}$ is a 3x3 neutrino flavor mixing unitary matrix retMaki . Here we assume that the neutrino wave-function have a plane-wave wavefunction retSwami and it propagates in vacuum from source $S(t_{S},\boldsymbol{x}_{S})$ to detector $D(t_{D},\boldsymbol{x}_{D})$, thus during the process of propagation, the phase accumulation of $\lvert\ \upsilon_{i}\rangle$ can be described by $\left\lvert\upsilon_{i}\left(t_{D},\boldsymbol{x}_{D}\right)\right\rangle=\exp\left(-\textup{i}\varphi_{i}\right)\left\lvert\upsilon_{i}\left(t_{S},\boldsymbol{x}_{S}\right)\right\rangle$, where $\varphi_{i}$ is the phase of oscillation. In flat spacetime, the phase is given by
@@
$\mathcal{P}_{\alpha\beta}\equiv\left\lvert\left\langle v_{\beta}\lvert v_{\alpha}(t_{D},\boldsymbol{x}_{D})\right\rangle\right\lvert^{2}=\sum_{i,j}U_{\beta i}U_{\beta j}^{*}U_{\alpha j}U_{\alpha i}^{*}exp[-\textup{i}(\varphi_{i}-\varphi_{j})],$ (6)
which means that when the initial and final flavor eigenstates are determined, the neutrino oscillation probability $\mathcal{P}_{\alpha\beta}$ depends only on the phase difference $\Delta\varphi_{ij}=\varphi_{i}-\varphi_{j}$, in the following section, we will focus on calculating the phase accumulation of neutrinos during their propagation in curved spacetime. Inspired by the work retBilenky:1978nj , we assume that all the mass eigenstates in a flavour eigenstate initially produced at the source $S(t_{s},x_{s})$ have equal energy or momentum.
@@ IV Phase of neutrino oscillation in the EHB gravity around global monopole field
Now let’s calculate relativistic neutrino oscillation phase described by the metric (Eq. (1)), for simplicity, the motion of neutrinos can be chosen to be confined on $\left(\theta=\frac{\pi}{2}\right)$ plane, thus the components of canonical momenta $p_{\mu}^{\left(h\right)}$ become
@@
$p_{r}^{\left(h\right)}=m_{h}g_{rr}\frac{dr}{ds}=m_{h}\frac{1+a}{A}\frac{dr}{ds},$
@@
$\varphi_{h}\approx\pm\int_{r_{s}}^{r_{D}}\frac{E_{h}}{A}\sqrt{1+a}\frac{Am_{h}^{2}}{2E_{h}^{2}}dr=\pm\sqrt{1+a}\frac{m_{h}^{2}}{2E_{0}}(r_{D}-r_{s}),$ (10)
@@ IV.2 Non-radial propagating neutrino phase in a null trajectory($d\phi\neq 0$)
The neutrino phase travelling along a null trajectory (the details see retChakrabarty:2023kld ) in terms of the non-radial case can be written as
@@
where $b$ is the impact parameter. We can see that the phase is not only affected by the geometry of spacetime, but also by the trajectory of the neutrino from the propagation source to the detector for the case of $b\neq 0$. Therefore, now we need to consider two cases discussing the phase:
(case 1) Neutrinos are produced at the source $S$, then propagate outwards non-radially without passing through dense celestial bodies until they reach the detector $D$. Since it is difficult to integrate Eq. (11) by replacing the metric coefficients, we rewrite it by using weak field approximation to get
@@
obviously, the above formula will reduce to that of the Schwarzschild case when $\overline{\mu}=a=0$ retNFO .
(case 2) Neutrinos are produced at the source $S$, propagate towards the gravitational lens, and finally reaches the detector $D$, here $r=r_{0}$ represents the closest distance to the lens during the path. Meanwhile we apply the weak field approximation, thus the phase transforms into
@@
$r_{0}\approx b\sqrt{1-\overline{\mu}}-\frac{M}{(1-\overline{\mu})}.$ (16)
@@ V Neutrino oscillation probability in the EHB gravity around global monopole field
As depicted in Fig. 1, neutrinos are produced in a same source and propagate outward along different paths. Due to the presence of a strong gravitational field (Schwarzschild-like black hole), the trajectories of some neutrinos deviate and converge at a common detection point $D$, resulting in the interference phenomenon, which is similar to how a convex lens (equivalent to a black hole) can converge the optical rays of the sun (equivalent to the propagation of neutrinos). Considering the fact that neutrinos traveling through different paths $p$ from S to D, so the neutrino flavour eigenstate need to be expressed as retChakrabarty
$|v_{\alpha}(t_{D},\boldsymbol{x}_{D})\rangle=N\sum_{i}U_{\alpha i}^{*}\sum_{p}\exp(-\textup{i}\varphi_{i}^{p})|v_{i}(t_{s},\boldsymbol{x}_{s})\rangle,$ (18)
where $\varphi_{i}^{p}$ is given by Eq.(17) and $N$ is a normalization factor. Considering the transition from an initial flavor eigenstate $\alpha$ at the source $S$ to a final flavor eigenstate $\beta$ at the detector $D$, the oscillation probability can be expressed as
$\mathcal{P}_{\alpha\beta}^{lens}\equiv\left|\left\langle v_{\beta}|v_{\alpha}\left(t_{D},\boldsymbol{x}_{D}\right)\right\rangle\right|^{2}=|N|^{2}\sum_{i,j}U_{\beta i}U_{\beta j}^{*}U_{\alpha j}U_{\alpha i}^{*}\sum_{p,q}\exp\left(-\textup{i}\Delta\varphi_{ij}^{pq}\right),$ (19)
@@
$|N|^{2}=\left(\sum_{i}\left|U_{\alpha i}\right|^{2}\sum_{p,q}\exp\left(-\textup{i}\Delta\varphi_{ii}^{pq}\right)\right)^{-1}.$ (20)
As mentioned before, when the initial and final flavor eigenstates are determined, the oscillation probability $\mathcal{P}_{\alpha\beta}^{lens}$ depends on the phase difference, thus in order to facilitate analysis, $\Delta\varphi_{ij}^{pq}$ can be represented by Eq.(17) as
@@
$A_{pq}=\frac{\sqrt{1+a}}{2E_{0}}\left(r_{D}+r_{s}\right)\left[1+\frac{2M}{\left(r_{D}+r_{s}\right)(1-\bar{\mu})}-\frac{(1-\bar{\mu})\sum b_{pq}^{2}}{4r_{D}r_{s}}\right],$ (22)
$B_{ij}=-\frac{(1-\bar{\mu})\sqrt{1+a}}{8E_{0}}\sum m_{ij}^{2}\left(\frac{1}{r_{D}}+\frac{1}{r_{S}}\right),$
@@
$\Delta m_{ij}^{2}=m_{i}^{2}-m_{j}^{2},$
Observing this form, we can see that if $p\neq q$, it is possible for the oscillation probability $\mathcal{P}_{\alpha\beta}^{lens}$ including the terms $\sum m_{ij}^{2}$ and $\Delta m_{ij}^{2}$, therefore, in principle, the observation of neutrinos in the lens could potentially detect the information about their absolute mass retChakrabarty .
@@
$\mathcal{P}_{e\mu}^{lens}$ $=|N|^{2}\sin^{2}2\gamma\bigg\{\frac{1}{2}\cos\bigl(\Delta b_{12}^{2}B_{11}\bigr)+\frac{1}{2}\cos\bigl(\Delta b_{12}^{2}B_{22}\bigr)+\sin^{2}\bigl(\frac{\Delta m_{21}^{2}A_{11}}{2}\bigr)$ (23)
$\quad+\sin^{2}\bigl(\frac{\Delta m_{21}^{2}A_{22}}{2}\bigr)-\cos\bigl(\Delta m_{21}^{2}A_{12}\bigr)\cos\bigl(\Delta b_{12}^{2}B_{12}\bigr)\bigg\},$
where $|N|^{2}=\frac{1}{2}\left[1+\cos^{2}\gamma\cos\left(\Delta b_{12}^{2}B_{11}\right)+\sin^{2}\gamma\cos\left(\Delta b_{12}^{2}B_{22}\right)\right]^{-1}$ and $\gamma$ is the mixing angle retEsteban:2018azc , we can see that $\mathcal{P}_{\alpha\beta}^{lens}$ depends on the neutrino mass ordering and for $\Delta m^{2}>0$ and $\Delta m^{2}<0$, it results in different results ($\gamma\neq\frac{\pi}{4}$).
VI Numerical results: two flavor case
We prefer to solve the path-dependent impact parameter $b_{p}$ in terms of the geometric quantities of the system. As shown in Fig. 1, it shows the description of lensing phenomena, in which the physical distance from the source to the lens is $r_{S}$ and the physical distance from the lens to the detector is $r_{D}$. For convenience, a new coordinate system $(x^{\prime},y^{\prime})$ is introduced by rotating the $(x,y)$ coordinate system through an angle $\Theta$ such that $x^{\prime}=x\cos{\Theta}+y\sin{\Theta}$ and $y^{\prime}=y\cos{\Theta}-x\sin{\Theta}$. Inspired by the work retCASA , the angle of deflection $\delta$ of neutrinos in the $(x^{\prime},y^{\prime})$ coordinate system can be expressed as